A mom is 30 years old and the daughter is 4. In how many years will the mother's age be triple her daughter's age is?
Answer by TimothyLamb(4379)

x = years in the future when mother's age is triple her daughter's age
---
30 + x = 3(4 + x)
30 + x = 12 + 3x
2x = 18
x = 9
---
answer:
in 9 years the mother's age will be triple her daughter's age
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php